# Scaling MLPs: A Tale of Inductive Bias

### v1

@@ Abstract
In this work we revisit the most fundamental building block in deep learning, the multi-layer perceptron (MLP), and study the limits of its performance on vision tasks. Empirical insights into MLPs are important for multiple reasons. (1) Given the recent narrative "less inductive bias is better", popularized due to transformers eclipsing convolutional models, it is natural to explore the limits of this hypothesis. To that end, MLPs offer an ideal test bed, being completely free of any inductive bias. (2) MLPs have almost exclusively been the main protagonist in the deep learning theory literature due to their mathematical simplicity, serving as a proxy to explain empirical phenomena observed for more complex architectures. Surprisingly, experimental datapoints for MLPs are very difficult to find in the literature, especially when coupled with large pre-training protocols. This discrepancy between practice and theory is worrying: Do MLPs reflect the empirical advances exhibited by practical models? Or do theorists need to rethink the role of MLPs as a proxy? We provide insights into both these aspects. We show that the performance of MLPs drastically improves with scale ($93\%$ on CIFAR10, $79\%$ on CIFAR100, $69\%$ on TinyImageNet), highlighting that lack of inductive bias can indeed be compensated. We observe that MLPs mimic the behaviour of their modern counterparts faithfully, with some components in the learning setting however surprisingly exhibiting stronger or unexpected behaviours. Due to their inherent computational efficiency, large pre-training experiments become more accessible for academic researchers. All of our experiments were run on a single GPU.
@@ 1 Introduction
In this work, we provide answers to question 1 and provide further evidence for hypothesis 2 and 3 by investigating how far we can push the empirical performance of models solely built from composing several MLP blocks. We give largely positive answers to question 1, observing that MLPs behave very similarly to their modern counterparts when subjected to scale, i.e. their performance increases predictably as a power law in parameter count and sample size, akin to Hestness et al., (2017, 2019); Kaplan et al., (2020); Zhai et al., (2022). In contrast to previous work however, we find that compute-optimal MLPs allocate their budget more strongly into sample size, highlighting again their lack of inductive bias. While regularization in the form of data augmentation is also helpful for CNNs, its role is significantly amplified for MLPs even at large sample sizes, leading to fatal degradation if turned off. We further investigate how the implicit bias of SGD affects performance, and we make a very counter-intuitive discovery: contrary to CNNs, we find that larger batch sizes generalize significantly better for MLPs. This result questions the validity of the proxy role that the MLP plays in theoretical works investigating the implicit bias of SGD. While, as expected, the scale employed in this work does not suffice for hypothesis 2, we provide strong evidence for 3, which we view as an important first step. We observe that enough scale indeed suffices to overcome the bad inductive bias present in MLPs, leading to surprisingly strong downstream performance, e.g. $\approx 93\%$ on CIFAR10, $\approx 79\%$ on CIFAR100 and $\approx 69\%$ on TinyImageNet. In summary, we make the following contributions:
@@ 2 Background
The MLP is a very general model and has no structure built into it to make it more suitable for specific tasks. A convolution on the other hand was designed specifically for vision with desirable characteristics incorporated into the model. A convolution can be viewed as a special case of an MLP, where the weight matrix $\bm{W}$ is very structured by being sparse and having shared entries, leading to spatially localized learning. This can be most easily illustrated in the case of convolving a $2\times 3\times 1$ image $\bm{x}$ with a $2\times 2$ filter $\bm{f}$ as the following matrix multiplication:
@@ 3 Architecture
with $\bm{W}\in\mathbb{R}^{m\times m}$. To embed the image $\bm{x}\in\mathbb{R}^{d\times d\times 3}$ we use a linear layer $\operatorname{emb}(\bm{x})=\bm{W}^{emb}\operatorname{vec}(\bm{x})$ with $\bm{W}^{emb}\in\mathbb{R}^{m\times 3d^{2}}$. Such an embedding layer is crucial since for high resolution images, $3d^{2}$ can be quite large and thus $m$ needs to be chosen smaller. We empirically find that such a network design is the minimal choice in order to guarantee successful training across all scales of parameter count and sample size.
@@
where $\bm{W}^{e}\in\mathbb{R}^{km\times m}$ expands the dimension to $km$ for $k\in\mathbb{N}$ and $\bm{W}^{(c)}\in\mathbb{R}^{m\times km}$ collapses it back to width $m$. For most experiments we set $k=4$. While the additions of skip connections and bottleneck layers to the architecture arguably add some amount of inductive bias, we believe that in comparison to modern architectures such enhancements remain negligible.
@@ 4.1 Setup
In this work, we solely focus on vision task as inductive bias is more readily understood in this setting. Moreover, most theoretical works focus on image classification tasks, making it thus a natural test bed to assess the performance of MLPs. We study the popular tasks CIFAR10, CIFAR100 (Krizhevsky,, 2009), STL10 (Coates et al.,, 2011), TinyImageNet (Le and Yang,, 2015), ImageNet1k for evaluation, as well as ImageNet21k (Deng et al.,, 2009) for pre-training. In order to limit the size of the embedding layer and the computational needs, we downscale all images to resolution $64\times 64\times 3$ (if needed) as done in Chrabaszcz et al., (2017). We center and normalize all the images as a pre-processing step. For data augmentations, we consider random flips and crops as well as MixUp (Zhang et al.,, 2018).
CIFAR10 CIFAR100 STL10 TinyImageNet ImageNet
B-6/Wi-1024 $69.9$±${\pm 0.1}$ $43.0$±${\pm 0.4}$ $51.5$±${\pm 0.1}$ $47.1$±${\pm 0.1}$ $15.2$±${\pm 0.2}$
B-6/Wi-1024 + DA $91.5$±${\pm 0.02}$ $76.4$±${\pm 0.2}$ $85.0$±${\pm 0.2}$ $62.7$±${\pm 0.1}$ $38.7$±${\pm 0.1}$
B-12/Wi-1024 + DA $93.6$±${\pm 0.03}$ $79.2$±${\pm 0.1}$ $88.9$±${\pm 0.1}$ $69.9$±${\pm 0.4}$ $42.5$±${\pm 0.2}$
@@ 4.2 Training from Scratch
We start the empirical exploration of MLPs by training them from scratch (i.e. without any extra data) on popular vision benchmarks. All models were trained with the LION optimizer (Chen et al.,, 2023) with a learning rate $\eta=5\mathrm{e}$-$5$. In order to combat overfitting we use strong label smoothing $\alpha=0.3$. We display the resulting test accuracies in Table 1. We first observe that the standard architecture of depth $L=6$ and width $m=1024$ without any data augmentation suffers from severe overfitting, leading to very suboptimal performance. Even when turning on data augmentation, it struggles to learn and performance gains are very modest. As observed in Lin et al., (2015), switching to the inverted bottleneck architecture improves performance across all datasets. Moreover, data augmentation as a regularizer really unfolds its full power, significantly pushing the performance by roughly $20\%$ across all tasks. Learning on the other hand significantly slows down with strong augmentations such as MixUp, enabling training for up to $5000$ epochs without suffering from overfitting. However, compared to simple modern baselines such as a ResNet18 (He et al.,, 2015), a large discrepancy in performance remains, highlighting the importance of inductive bias in the small sample regime. We remark that ViTs and MLP-Mixers as well exhibit more learning difficulties if the dataset size is small (Dosovitskiy et al.,, 2021; Tolstikhin et al.,, 2021). We provide more ablation studies in Appendix A.2.
@@ 4.3 Transfer Learning
In this section, we aim to analyze how transferable features learnt by MLPs are across different vision tasks. Transferability is one of the hallmark characteristics of modern deep learning, enabling practitioners to fine-tune large models on their specific dataset, leading to superior performance. We are, to the best of our knowledge, the first to measure transferability of MLPs, which is crucial to assess in order to build a theoretical understanding of the process. In this section, we focus on the inverted bottleneck MLP as it generalizes better and is easier to optimize. We restrict to $k=4$ for the expansion factor and denote by B-L/Wi-m a network with $L$ blocks and width $m$. For pre-training we use ImageNet21k, the largest publicly available image dataset with annotated classes. After preprocessing the dataset following Ridnik et al., (2021), it consists of roughly 12 million images and 11 thousand classes. We then pre-train the MLP with the cross-entropy loss for 400 epochs, employing label smoothing and the LION optimizer. To guarantee fast data loading we rely on the FFCV framework (Leclerc et al.,, 2023) for all experiments.
@@
In order to measure transferability of the learnt features we fine-tune the network on the new task. We also study training a linear layer on top of the embeddings but defer those results to Appendix A.3. We again explore the effects of data augmentation during the pre-training stage. For fine-tuning we use SGD with momentum with a learning rate of $\eta_{\text{head}}=0.01$ for the head and $\eta_{\text{body}}=0.001$ for the encoder for $50$ epochs. We upscale CIFAR images to resolution $64\times 64\times 3$ at fine-tuning time to guarantee compability. We display the fine-tuning results in Table 2. For visualizations of the learnt features, we refer the interested reader to Appendix 11. We again observe that using data augmentation during the pre-training phase is essential to successful training, boosting performance up to $30\%$ in case of CIFAR100. Surprisingly, the learnt features are highly transferable, improving the performances reported previously in Table 1 dramatically. While of course pre-trained on a large quantity of data, we nevertheless want to highlight that such an MLP becomes competitive with a ResNet18 trained from scratch for all the datasets, except for ImageNet1k where performance falls surprisingly short. We hypothesize that MLPs struggle with the more fine-grained distinctions between classes, in combination with the reduced resolution of the images. Overall however, these results underline that a bad inductive bias as exhibited by an MLP can indeed be overcome if subjected to enough scale. For theory, the results are double-edged; while MLPs prove to be a good proxy to understand transfer learning, data augmentation and the inverted bottleneck structure seem to be an essential component to the success. Both these characteristics on the other hand remain rather understudied in theoretical works.
@@
The role of data augmentation is very pronounced for MLPs, largely since it provides indirect inductive bias to the model. Remarkably, a model pre-trained on $12$ million examples without data augmentation shows inferior performance on CIFAR10 compared to a network trained from scratch with augmentations turned on. This emphasizes that augmentations go beyond merely leading to a bigger dataset but provide the model with useful invariances. We investigate the learnt weights in-depth in Appendix 11, showing that very evidently, more localized features are learnt if data augmentation is employed.
@@ 4.4 Scaling Laws
One of the key mysteries in deep learning is that networks tend to improve in terms of generalization when compute, in the form of parameter count and dataset size, is scaled up. Recently it has been observed in several works that the benefits of scale are highly predictable, i.e. generalization performance exhibits a power-law structure when plotted against compute measured in FLOPS (Hestness et al.,, 2017, 2019; Kaplan et al.,, 2020; Zhai et al.,, 2022). This predictability has even been leveraged to estimate the optimal model before training (Hoffmann et al.,, 2022; OpenAI,, 2023). In order to understand this important characteristic of deep learning theoretically, it is important to analyze whether MLPs exhibit similar properties.
@@
where $a,b,E_{\infty}\in\mathbb{R}_{+}$ and $\alpha>0$ is the scaling coefficient determining the rate of decay. $E_{\infty}$ denotes the irreducible error, i.e. even if infinite compute were employed, the performance remains imperfect. The test error can be measured upstream (i.e. on the pre-training task) or downstream when fine-tuning on a different task. We investigate various pre-training schemes with different number of examples, parameter counts and training times. We subsample ImageNet21k proportionally across classes and pre-train variously sized inverted bottleneck MLPs. We summarize the configurations in Table 5. We then measure test error on the downstream task of CIFAR100 in Fig. 1 as well as CIFAR10 and ImageNet1k in Fig. 6 by linearly transferring the learnt features. The plotting style is inspired by Zhai et al., (2022). Each point in the curve is the downstream performance of an MLP, where the color of the point indicates the model type (blue denotes smaller and red larger models) and the size of the point indicates the number of pre-training examples. Points connected by a line indicates longer training times where $T\in\{50,100,200,400,800\}$ is measured in epochs. In all experiments, we employ data augmentation for pre-training. We observe that the compute-optimal performance of MLPs strongly exhibits characteristics of a power-law with coefficients $\alpha\in\{0.12,0.25,0.35\}$. This is very encouraging for future theoretical work, showing that MLPs indeed mirror the scaling behaviour of modern models.

### v3

@@ Abstract
In this work we revisit the most fundamental building block in deep learning, the multi-layer perceptron (MLP), and study the limits of its performance on vision tasks. Empirical insights into MLPs are important for multiple reasons. (1) Given the recent narrative "less inductive bias is better", popularized due to transformers eclipsing convolutional models, it is natural to explore the limits of this hypothesis. To that end, MLPs offer an ideal test bed, as they lack any vision-specific inductive bias. (2) MLPs have almost exclusively been the main protagonist in the deep learning theory literature due to their mathematical simplicity, serving as a proxy to explain empirical phenomena observed for more complex architectures. Surprisingly, experimental datapoints for MLPs are very difficult to find in the literature, especially when coupled with large pre-training protocols. This discrepancy between practice and theory is worrying: Do MLPs reflect the empirical advances exhibited by practical models? Or do theorists need to rethink the role of MLPs as a proxy? We provide insights into both these aspects. We show that the performance of MLPs drastically improves with scale ($95\%$ on CIFAR10, $82\%$ on CIFAR100, $58\%$ on ImageNet ReaL), highlighting that lack of inductive bias can indeed be compensated. We observe that MLPs mimic the behaviour of their modern counterparts faithfully, with some components in the learning setting however exhibiting stronger or unexpected behaviours. Due to their inherent computational efficiency, large pre-training experiments become more accessible for academic researchers. All of our experiments were run on a single GPU.
@@ 1 Introduction
In this work, we provide answers to question 1 and provide further evidence for hypothesis 2 and 3 by investigating how far we can push the empirical performance of models solely built from composing several MLP blocks. We give largely positive answers to question 1, observing that MLPs behave very similarly to their modern counterparts when subjected to scale, i.e. their performance increases predictably as a power law in parameter count and sample size, akin to Hestness et al., (2017, 2019); Kaplan et al., (2020); Zhai et al., (2022) (see e.g. Fig. 1). In contrast to previous work however, we find that compute-optimal MLPs allocate their budget more strongly into sample size, highlighting again their small inductive bias. While regularization in the form of data augmentation is also helpful for CNNs, its role is significantly amplified for MLPs even at large sample sizes, leading to fatal degradation if turned off. We further investigate how the implicit bias of SGD affects performance, and we make a very counter-intuitive discovery: contrary to CNNs, we find that larger batch sizes generalize significantly better for MLPs. This result questions the validity of the proxy role that the MLP plays in theoretical works investigating the implicit bias of SGD. While, as expected, the scale employed in this work does not suffice for hypothesis 2, we provide strong evidence for 3, which we view as an important first step. We observe that enough scale indeed suffices to overcome the bad inductive bias present in MLPs, leading to surprisingly strong downstream performance, e.g. $\approx 95\%$ on CIFAR10, $\approx 82\%$ on CIFAR100 and $\approx 58\%$ on ImageNet ReaL. In summary, we make the following contributions:
@@ 2 Background
We want to highlight that MLPs of course are not completely free of inductive bias, in the sense that they encourage learning a hierarchical feature structure. On the other hand, there is no vision-specific inductive bias present in MLPs, which is the main setting we investigate here. We refer to Battaglia et al., (2018) for a more in-depth treatment of inductive bias.
@@
The MLP is a very general model and has no structure built into it to make it more suitable for vision tasks. A convolution on the other hand was designed specifically for vision with desirable characteristics incorporated into the model. A convolution can be viewed as a special case of an MLP, where the weight matrix $\bm{W}$ is very structured by being sparse and having shared entries, leading to spatially localized learning. This can be most easily illustrated in the case of convolving a $2\times 3\times 1$ image $\bm{x}$ with a $2\times 2$ filter $\bm{f}$ as the following matrix multiplication:
@@
Patchifiying.
As highlighted above, ViTs and Mixers largely obtain their inductive biases through breaking the images into patches. This choice seems to be beneficial even for architectures that already possess a strong inductive bias, such as the ConvMixer (Trockman and Kolter,, 2022), where convolutions are performed on individual patches. The very recent Metaformer (Yu et al.,, 2022) further shows that even a simple spatial pooling instead of attention can lead to strong performance if the image is patchified. While the success of this mechanism certainly warrants further investigation, in this work we decided to deliberately focus on MLPs as they specifically lack this type of bias.
@@ 3 Architecture
with $\bm{W}\in\mathbb{R}^{m\times m}$. To embed the image $\bm{x}\in\mathbb{R}^{d\times d\times 3}$ we use a linear layer $\operatorname{emb}(\bm{x})=\bm{W}^{emb}\operatorname{vec}(\bm{x})$ with $\bm{W}^{emb}\in\mathbb{R}^{m\times 3d^{2}}$. Such an embedding layer is crucial since for high resolution images, $3d^{2}$ can be quite large and thus $m$ needs to be chosen smaller. We empirically find that such a network design is the minimal choice in order to guarantee successful training across all scales of parameter count and sample size. We will use the short cut S-MLP to denote such an architecture.
@@
where $\bm{W}^{e}\in\mathbb{R}^{km\times m}$ expands the dimension to $km$ for $k\in\mathbb{N}$ and $\bm{W}^{(c)}\in\mathbb{R}^{m\times km}$ collapses it back to width $m$. For most experiments we set $k=4$. While the additions of skip connections and bottleneck layers to the architecture arguably add some amount of inductive bias, we believe that in comparison to modern architectures such enhancements remain negligible. We will denote this variant by B-MLP.
@@ 4.1 Setup
In this work, we solely focus on vision tasks as inductive bias is more readily understood in this setting. Moreover, most theoretical works focus on image classification tasks, making it thus a natural test bed to assess the performance of MLPs. We study the popular tasks CIFAR10, CIFAR100 (Krizhevsky,, 2009), STL10 (Coates et al.,, 2011), TinyImageNet (Le and Yang,, 2015), ImageNet1k for evaluation, as well as ImageNet21k (Deng et al.,, 2009) for pre-training. In order to limit the size of the embedding layer and the computational needs, we downscale all images to resolution $64\times 64\times 3$ (if needed) as done in Chrabaszcz et al., (2017). We center and normalize all the images as a pre-processing step. For data augmentations, we consider random flips and crops as well as MixUp (Zhang et al.,, 2018).
CIFAR10 CIFAR100 STL10 Tiny-IN IN ReaL
B-6/Wi-1024 $69.9$±${\pm 0.1}$ $43.0$±${\pm 0.4}$ $51.5$±${\pm 0.1}$ $47.1$±${\pm 0.1}$ $15.2$±${\pm 0.2}$ ${\pm 0.2}$
B-6/Wi-1024 + DA $91.5$±${\pm 0.02}$ $76.4$±${\pm 0.2}$ $85.0$±${\pm 0.2}$ $62.7$±${\pm 0.1}$ $38.7$±${\pm 0.1}$ $47.0$±${\pm 0.15}$
B-12/Wi-1024 + DA $94.2$±${\pm 0.05}$ $80.0$±${\pm 0.05}$ $89.9$±${\pm 0.1}$ $69.9$±${\pm 0.4}$ $43.3$±${\pm 0.06}$ $48.6$±${\pm 0.2}$
B-12/Wi-1024 + DA + TTA $95.5$±${\pm 0.05}$ $82.6$±${\pm 0.2}$ $92.2$±${\pm 0.05}$ $73.1$±${\pm 0.5}$ $51.5$±${\pm 0.1}$ $57.9$±${\pm 0.1}$
@@ 4.2 Training from Scratch
We start the empirical exploration of MLPs by training them from scratch (i.e. without any extra data) on popular vision benchmarks. All models were trained with the LION optimizer (Chen et al.,, 2023) with a learning rate $\eta=5\mathrm{e}$-$5$. In order to combat overfitting we use strong label smoothing $\alpha=0.3$. We display the resulting test accuracies in Table 1. We observe that both the standard architecture and the bottleneck without any data augmentation suffer from overfitting, leading to suboptimal performance. When turning it on, data augmentation as a regularizer however really unfolds its full power, significantly pushing the performance by roughly $20\%$ across all tasks. As observed in Lin et al., (2015), the inverted bottleneck architecture leads to an improvement in performance across all datasets. Learning on the other hand significantly slows down with strong augmentations such as MixUp, enabling training for up to $5000$ epochs without suffering from overfitting. However, compared to simple modern baselines such as a ResNet18 (He et al.,, 2015), a large discrepancy in performance remains, highlighting the importance of inductive bias in the small sample regime. We remark that ViTs and MLP-Mixers as well exhibit more learning difficulties if the dataset size is small (Dosovitskiy et al.,, 2021; Tolstikhin et al.,, 2021). We provide more ablation studies in Appendix A.2.
@@ 4.3 Transfer Learning
In this section, we aim to analyze how transferable features learnt by MLPs are across different vision tasks. Transferability is one of the hallmark characteristics of modern deep learning, enabling practitioners to fine-tune large models on their specific dataset, leading to superior performance. We are, to the best of our knowledge, the first to measure transferability of MLPs, which is crucial to assess in order to build a theoretical understanding of the process. In this section, we focus on the inverted bottleneck MLP as it generalizes better and is easier to optimize. We provide the dual results for the standard MLP in Appendix B.1. We restrict to $k=4$ for the expansion factor and denote by B-L/Wi-m a network with $L$ blocks and width $m$. For pre-training we use ImageNet21k, the largest publicly available image dataset with annotated classes. After preprocessing the dataset following Ridnik et al., (2021), it consists of roughly 12 million images and 11 thousand classes. We then pre-train the MLP with the cross-entropy loss for $800$ epochs, employing label smoothing and the LION optimizer. To guarantee fast data loading we rely on the FFCV framework (Leclerc et al.,, 2023) for all experiments.
@@
In order to measure transferability of the learnt features we fine-tune the network on the new task. We also study training a linear layer on top of the embeddings but defer those results to Appendix A.3. We again explore the effects of data augmentation during the pre-training stage. For fine-tuning we use SGD with momentum with a learning rate of $\eta_{\text{head}}=0.01$ for the head and $\eta_{\text{body}}=0.001$ for the encoder for $50$ epochs. We upscale CIFAR images to resolution $64\times 64\times 3$ at fine-tuning time to guarantee compatibility. We display the fine-tuning results in Table 2. For visualizations of the learnt features, we refer the interested reader to Appendix 12. We again observe that using data augmentation during the pre-training phase is essential to successful training, boosting performance up to $30\%$ in case of CIFAR100. Surprisingly, the learnt features are highly transferable, improving the performances reported previously in Table 1 dramatically. While of course pre-trained on a large quantity of data, we nevertheless want to highlight that such an MLP becomes competitive with a ResNet18 trained from scratch for all the datasets, except for ImageNet1k where performance falls surprisingly short. We hypothesize that MLPs struggle with the more fine-grained distinctions between classes, in combination with the reduced resolution of the images.
Test-Time Augmentations.
For ImageNet1k we further notice that objects tend to not be centered, in contrast to datasets like CIFAR10. We suspect that this might lead to the comparatively weaker performance. To test this, we leverage test-time augmentations (TTA). As introduced by Krizhevsky et al., (2012), for each test image, we produce a fixed number of $100$ random crops and use the averaged logits for prediction. We observe significant improvements across all datasets, especially for ImageNet we obtain an increase of roughly $8\%$. This indeed indicates that MLPs struggle to localize the object of interest, especially for the more complicated ImageNet1k task. Using a large number of crops alleviates this problem to some degree. This also explains why the gains on tasks like CIFAR10 are smaller as the objects there usually are perfectly centered.
ReaL accuary.
As observed in (Beyer et al.,, 2020), the ImageNet labels do not capture that a single image might contain multiple objects of distinct classes. ImageNet accuracy can thus be misleading in the sense that model classes such as convolutional networks might have implicitly adapted to the particular labeling strategy due to the repeated benchmarking on the same validation set. MLPs most likely lack such an implicit adaptation as this work is to our knowledge the first to evaluate them on ImageNet1k. To address this, Beyer et al., (2020) introduced a novel set of validation labels that better capture the multi-label nature, where a prediction is deemed correct if it matches one of the categories present in the image. We observe further very significant improvements of $\approx 7\%$ when employing ImageNet ReaL.
Overall, these results underline that a bad inductive bias as exhibited by an MLP can indeed be overcome if subjected to enough scale. For theory, the results are double-edged; while MLPs prove to be a good proxy to understand transfer learning, data augmentation proves to be a crucial component. Also test-time augmentations significantly boost performance. Both these components on the other hand remain rather understudied in theoretical works.
@@
The role of data augmentation is very pronounced for MLPs, largely since it provides indirect inductive bias to the model. Remarkably, a model pre-trained on $12$ million examples without data augmentation shows inferior performance on CIFAR10 compared to a network trained from scratch with augmentations turned on. This emphasizes that augmentations go beyond merely leading to a bigger dataset but provide the model with useful invariances. We investigate the learnt weights in-depth in Appendix 12, showing that very evidently, more localized features are learnt if data augmentation is employed. The power of augmentations has already been demonstrated previously through the advent of self-supervised learning (Grill et al.,, 2020; Caron et al.,, 2021; Chen et al.,, 2020). Even when training on purely random labels, it still provides a powerful learning signal (Anagnostidis et al.,, 2023).
@@ 4.4 Scaling Laws
One of the key mysteries in deep learning is that networks tend to improve in terms of generalization when compute, in the form of parameter count and dataset size, is scaled up. Recently it has been observed in several works that the benefits of scale are highly predictable, i.e. generalization performance exhibits a power-law structure when plotted against compute measured in FLOPS (Rosenfeld et al.,, 2020; Hestness et al.,, 2017, 2019; Kaplan et al.,, 2020; Zhai et al.,, 2022). The functional form has recently been further refined (Caballero et al.,, 2023). The predictable nature of test performance has even been leveraged to estimate the optimal model before training (Hoffmann et al.,, 2022; OpenAI,, 2023). In order to understand this important characteristic of deep learning theoretically, it is important to analyze whether MLPs exhibit similar properties.
@@
where $a,b,E_{\infty}\in\mathbb{R}_{+}$ and $\alpha>0$ is the scaling coefficient determining the rate of decay. $E_{\infty}$ denotes the irreducible error, i.e. even if infinite compute were employed, the performance remains imperfect. The test error can be measured upstream (i.e. on the pre-training task) or downstream when fine-tuning on a different task. We investigate various pre-training schemes with different number of examples, parameter counts and training times. We subsample ImageNet21k proportionally across classes and pre-train variously sized inverted bottleneck MLPs. We summarize the configurations in Table 5. We then measure test error on the downstream task of CIFAR100 in Fig. 1 as well as CIFAR10 and ImageNet1k in Fig. 6 by linearly transferring the learnt features (without test-time augmentations). The plotting style is inspired by Zhai et al., (2022). Each point in the curve is the downstream performance of an MLP, where the color of the point indicates the model type (blue denotes smaller and red larger models) and the size of the point indicates the number of pre-training examples. Points connected by a line indicates longer training times where $T\in\{50,100,200,400,800\}$ is measured in epochs. In all experiments, we employ data augmentation for pre-training. We observe that the compute-optimal performance of MLPs strongly exhibits characteristics of a power-law with coefficients $\alpha\in\{0.12,0.25,0.35\}$. This is very encouraging for future theoretical work, showing that MLPs indeed mirror the scaling behaviour of modern models. We provide the dual results for the standard MLPs in Appendix B.2, noting that they exhibit essentially the same scaling behaviour, albeit with a slightly weaker slope and intercept.